OpenStudy (anonymous):

Population of a town can be modeled by P=2|t-8|+4, where t=0 represents 1990. During which two years does the town have a population of 8000?

OpenStudy (anonymous):

Plug in 8,000 for P and solve for t. There will be two answers because there is an absolute value. Keep in mind that t only equals the NUMBER of years after 1990; to find the YEAR, you need to add those numbers to 1990.

OpenStudy (anonymous):

put \[8000=2|t-8|+4\] \[7996=2|t-8|\] \[3998=|t-8|\] \[t-8=3998\] \[t=4006\] so one answer is \[1990+4006\] what kind of weird question is this? other one is \[t-8=-3998\] so \[t=-3990\] etc

OpenStudy (anonymous):

^ Lol, satellite73 is correct; there is something wrong with the population model..